Find values for m and b in the following system so that the solution to the system is (-3, 4).
5x + 7y = b
mx + y = 22
Please help.

Answer by checkley75(3666)

5x+7y=b
5*-3+7*4=b
-15+28=b
b=13 answer.
mx+y=22
m*-3+4=22
-3m=22-4
-3m=18
m=18/-3
m=-6 answer.